Samuel Hocevar

6 bis passage Dareau 75 014 Paris

Phone: (+33) 681 122 062

E-mail: sam@hocevar.net

Born on 5th August 1978 in Forbach (Moselle) French nationality French driving licence

Languages

French  fluent (mother tongue), good grammar and spelling

English fluent (TOEFL score: 647 in 1997), numerous stays in England, can tell the difference between it’s and its

German  good knowledge and practice
Spanish  good knowledge

Professional experience

2009–present Senior R&D Engineer at DONTNOD Entertainment (C, C++, Unreal Engine, fluid simulations, PS3, Xbox 360, PowePC, SPU, Cell, AltiVec, SSE3)

2006–2009 Director of R&D at [redacted], a 700-employee company (surveillance systems, video streaming solutions, intelligent vision)

2003–2006 (not de- fended yet)

SNCF - Direction de l’Innovation et de la Recherche, ENST - Laboratoire Informatique et Réseaux: Computer Science Industrial PhD ("Generation of realistic railroad simulations from video images") (image processing, computer vision, 3D reconstruction, video coding, HDTV, C, Unix)

2002 SNCF - Direction de la Recherche et de la Technologie, unité Nouvelles Technologies de l’Information et de la Communication (Paris, France), 8 months: work on a showpiece for at-seat multimedia services and video streaming research (C, C++, Unix, Win32, WinCE, MPEG, MPEG2, MPEG4, WiFi)

1998–2002 VideoLAN, video streaming solution and cross-platform DVD player: major project contributor for several years (C, C++, Unix, Win32, MPEG, MPEG2, DVD)

1999 ClaraNet France (Paris, France), 7 months: systems team - Unix administration, nextwork tools development (C, PHP, MySQL)

Training

2003–2006 (unfinished) École Nationale Supérieure des Télécomunications (ENST), Paris: Computer Science Industrial PhD

1997–2002 École Centrale of Paris: engineer’s diploma, electromechanical systems and applied mechatronics speciality, project management specialisation, graduated with honours (mention bien)

1995–1997  Lycée Fabert, Metz (France, Moselle): preparatory classes

Personal knowledge

Operating systems - GNU/Linux (administration well-mastered, part of the Debian project)

- 3rd generation console systems (PS3 & Xbox 360), iOS, Android

- Solaris, BSD Unix, Mac OS X, GNU/Hurd, Windows

Networks  - LAN (Ethernet, ARP, RARP, Token Ring), ATM

- Cisco IOS, SNMP aministration and monitoring

- low-level Internet: IP, TCP, UDP, ICMP, IGMP

- high-level Internet: HTTP, NNTP, SMTP, DNS, DHCP, NFS

Administration - 9 years of Unix administration experience, excellent knowledge of the common tools

- examples of well-known server setups: INN, Apache, BIND, Postfix

Programming - low-level: x86 and amd64 assembly (MMX, SSE, SSE3, excellent knowledge), PowerPC & Cell assembly (AltiVec, SPU), NEC/Saturn assembly (mastered), Z80 and m68k assembly, reverse-engineering

- high-level: C & C++ (excellent knowledge, several personal and team projects of more than 5,000 lines), C#, Java, Scheme (good knowledge), bases in Pascal and Caml

- libraries: Xlib, GTK+/Gnome, SDL

- scripting languages: POSIX shell, PHP, sed and awk well mastered, good knowledge of Python, Perl

Other  - databases: MySQL, PostgreSQL

- markup languages: LaTeX2e, SGML, HTML, XHTML, CSS

- good knowledge of the MPEG and MPEG2 standards (especially the system layer), excellent knowledge of the DVD format

- excellent image processing skills (various projects, from robot guiding to captcha decoding)

Miscellanea

Sports  - practical experience of judo for 8 years, soccer, badminton, coneball

- chess competitions (current ELO: 1860)

Culture and leisure - drawing and painting, calligraphy, reading, cinema, XXth century artistic movements, sabotage of oppressive structures, cooking

- active contributor of the Wikipedia project (French and English versions) since 2004

Computers - part of the Debian project since 1999 (Project Leader 2007-2008), various activities amongst the OpenSource community, mostly development